James has two planks of wood. One of them is 16m less than 3 times the length of the other one . The total length of the two planks is 8 metres long.
If one of the planks of wood has a length of x metres long. Write an algebraic expression of the other plank

Answer by ikleyn(52944)

.
Second plank: x meters long.

First plank: 3x - 16 meters long.

Both together are x + (3x-16) = 4x - 16.


Your equation is 

4x - 16 = 8.


Your solution is 

4x = 8 + 16 = 24  ====>  x = 6.


Answer.  Second plank is 3 m.  First plank is 3x-16 = 3*6 - 16 = 2m.

Solved.